A gym owner is considering opening a location on the other side of town. The new facility will cost $1.38 million and will be depreciated on a straight-line basis over a 20-year period. The new gym is expected to generate $541,000 in annual sales. Variable costs are 52 percent of sales, the annual fixed costs are $86,900, and the tax rate is 34 percent. What is the operating cash flow?
Answer #1

Sales $ 541,000 $ =541000*0.52+86900 =1380000/20 Less: Costs Depreciation EBIT Less: Tax @ 34% Net income Add: Depreciation $

Please rate thumbs up